OpenStudy (anonymous):

Differentiate function r(t)= 5t^(-3/5)

OpenStudy (anonymous):

wouldn't that be -3/t^8/5?

OpenStudy (anonymous):

Im pretty sure

OpenStudy (anonymous):

off by a minus sign in your exponent

OpenStudy (anonymous):

should be \[-3t^{-\frac{8}{5}}\]

OpenStudy (anonymous):

so its not -3 divided by t^-8/5?

OpenStudy (anonymous):

it is actually. d/dx 5x^(-3/5) = (-3/5)(5)=-3^-8/5

OpenStudy (anonymous):